What is the volume of a cube with one side 3.5cm long?
Answer by rapaljer(4667)

If the side of the cube is x, then the volume is V=x^3.

In this case, V=3.5^3 = 42.875 cubic cm.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida